### v2

# Distributional Balancing for Causal Inference: A Unified Framework via Characteristic Function Distance

Abstract
Weighting methods are essential tools for estimating causal effects in observational studies, with the goal of balancing pre-treatment covariates across treatment groups. Traditional approaches pursue this objective indirectly, for example, via inverse propensity score weighting or by matching a finite number of covariate moments, and therefore do not guarantee balance of the full joint covariate distributions. Recently, distributional balancing methods have emerged as robust, nonparametric alternatives that directly target alignment of entire covariate distributions, but they lack a unified framework, formal theoretical guarantees, and valid inferential procedures. We introduce a unified framework for nonparametric distributional balancing based on the characteristic function distance (CFD) and show that widely used discrepancy measures, including the maximum mean discrepancy and energy distance, arise as special cases. Our theoretical analysis establishes conditions under which the resulting CFD-based weighting estimator achieves $\sqrt{n}$-consistency. Since the standard bootstrap may fail for this estimator, we propose subsampling as a valid alternative for inference. We further extend our approach to an instrumental variable setting to address potential unmeasured confounding. Finally, we evaluate the performance of our method through simulation studies and a real-world application, where the proposed estimator performs well and exhibits results consistent with our theoretical predictions.
Keywords: Energy distance, Local average treatment effect, Maximum mean discrepancy, Reproducing kernel Hilbert space, Quadratic programming, Subsampling
1 Introduction
A primary objective of causal inference is to estimate causal effects in both randomized experiments and observational studies. In well-executed randomized experiments, treatment assignment is independent of potential outcomes, allowing differences in observed outcomes to be interpreted as causal effects. By contrast, in observational studies, treatment assignment may be confounded with potential outcomes, making naive comparisons of outcomes cannot be interpreted as causal effects and require appropriate adjustments to address confounding bias. A popular approach to addressing confounding in observational studies is weighting, which weights observations to create a pseudo-population in which treated and control groups are more comparable, thereby reducing bias in causal effect estimates; see Section 2.2 for a review of weighting estimators.
Among these methods, inverse probability weighting (IPW; Horvitz and Thompson, 1952) is the most widely used (Austin and Stuart, 2015; Hirano et al., 2003; Lee et al., 2010; Robins et al., 2000; Robins and Rotnitzky, 1995). In general, IPW methods weight each observation by the reciprocal of its propensity score (Rosenbaum and Rubin, 1983), defined as the probability of receiving treatment conditional on observed covariates. However, for their validity, IPW methods typically rely on the correct specification of the propensity score model, as misspecification can introduce substantial bias. Consequently, recent developments in weighting have focused on (near-)balancing covariate moments to improve robustness (Chan et al., 2015; Cohn et al., 2023; Hainmueller, 2012; Imai and Ratkovic, 2014; Soriano et al., 2023; Zubizarreta, 2015). Nevertheless, challenges remain in deciding which moments to balance, and balancing a finite set of moments does not guarantee elimination of confounding beyond those moments. These limitations of IPW and moment-based balancing methods can reduce their reliability and practical applicability in real-world settings; see Sections 6 and 7 for simulation studies and real-world applications illustrating these issues.
To address these limitations, recent studies have proposed distributional balancing methods (Ben-Michael et al., 2021; Bruns-Smith et al., 2025; Chen et al., 2024; Clivio et al., 2024; Huling et al., 2024; Huling and Mak, 2024; Kallus, 2018, 2020; Kong et al., 2023; Wang et al., 2023) which typically proceed in two steps. First, one selects a discrepancy measure to compare full covariate distributions, rather than finite-dimensional summaries such as moments, by comparing the weighted covariate distribution within each treatment group to the unweighted covariate distribution in the overall sample. Second, balancing weights are obtained by minimizing this discrepancy via an optimization procedure (see Section 3.2 for an example). Unlike IPW and moment-based methods, whose validity hinges on correct specification of the propensity score and the sufficiency of a finite set of moments for removing confounding, distributional balancing directly targets imbalance in the full covariate distributions under weaker assumptions. In practice, such weights can improve empirical performance by achieving balance for arbitrary functions of the covariates without committing to a particular propensity score model or moment conditions that are often unrealistic.
Despite the growing literature, distributional balancing methods remain underdeveloped. First, most existing methods are designed in isolation, tailored to specific distance metrics, and lack a unified framework that accommodates a broad class of metrics. Second, formal theoretical analyses of the resulting weighting estimators are often absent. In particular, it remains unclear whether standard asymptotic properties, such as $\sqrt{n}$-consistency, hold for estimators derived from a given distance metric, and under what conditions these properties are satisfied. Moreover, formal procedures for conducting valid statistical inference on causal effects based on distributional balancing weights remain largely unexplored and often rely naively on bootstrap methods.
In this paper, we develop a novel nonparametric distributional balancing method that addresses multiple limitations of existing approaches. The main contributions of our work can be summarized as follows:
  • 1.
    We propose a method based on a characteristic function distance (CFD; Ansari et al., 2020; Heathcote, 1977; Chwialkowski et al., 2015), which generalizes several popular distributional balancing methods across a broad class of distance functions. Notably, the construction of the CFD depends on the choice of a density function $\omega$, with certain selections of $\omega$ corresponding directly to established methods. In particular, previous approaches based on the energy distance (Huling and Mak, 2024) and the maximum mean discrepancy (Chen et al., 2024) are special cases of our method when $\omega$ is appropriately specified; see Section 3.3 for a detailed discussion of these relationships. Importantly, our framework is more general than previous methods, as $\omega$ can be chosen flexibly beyond the specific forms required by earlier approaches, provided it satisfies mild regularity conditions.
  • 2.
    We clarify how the choice of $\omega$ affects the feasibility of achieving $\sqrt{n}$-consistency for the resulting weighting estimator. Roughly speaking, the tail behavior of $\omega$ determines the smoothness requirements on the outcome regression (i.e., the treatment-specific covariate-outcome relationship). Therefore, we formally establish that distributional balancing with different CFDs necessitates different assumptions on the outcome regression, highlighting that not all choices of $\omega$ guarantee $\sqrt{n}$-consistent estimation for a given dataset; see Section 4 for the details.
  • 3.
    We show that bootstrap-based inference may fail to produce reliable results because weighting estimators based on distributional balancing weights may not satisfy the conditions necessary for valid bootstrap inference; see Section 6 for empirical evidence from simulation studies. As an alternative, we propose using subsampling (Politis et al., 1999), which remains valid provided that the weighting estimator is $\sqrt{n}$-consistent.
  • 4.
    In Section 5, we extend our approach to relax the no unmeasured confounding assumption. Specifically, within the instrumental variable (IV) framework of Angrist et al. (1996), we show that our distributional balancing method can be directly applied to estimate the local average treatment effect via the Wald estimand in equation (20). Unlike other regression-based approaches, our weighting estimator based on distributional balancing weights is a one-shot method that does not require estimating multiple nuisance components. We further demonstrate the broader applicability of our method through simulation studies and a real-world application under this IV setting. The results indicate that the CFD-based weighting estimator performs consistently with our theoretical predictions; see Sections 6 and 7 for details.
2 Setup & Review
2.1 Setup
Consider an independent and identically distributed (i.i.d.) sample $\{(Y_{i},Z_{i},\mathbf{X}_{i})\}_{i=1}^{n}$ of size $n$ generated from a population, where, for unit $i\in\{1,\ldots,n\}$, $Y_{i}\in\mathbb{R}$ denotes the observed outcome, $Z_{i}\in\{0,1\}$ denotes an indicator of whether the unit receives treatment $(Z_{i}=1)$ or control $(Z_{i}=0)$, and $\mathbf{X}_{i}\in\mathcal{X}\subset\mathbb{R}^{d}$ denotes a $d$-dimensional vector of pre-treatment covariates. We use $F$ and $\mathcal{X}$ to denote the distribution of $\mathbf{X}$ and its support, respectively. Let $Y_{i}(z)\in\mathbb{R}$ be the potential outcome of unit $i\in\{1,\ldots,n\}$, had possibly contrary to fact, the treatment been set to $Z_{i}=z\in\{0,1\}$. Let $n_{1}=\sum_{i=1}^{n}Z_{i}$ and $n_{0}=\sum_{i=1}^{n}(1-Z_{i})$ denote the number of units received treatment and control, respectively. In addition, we define the empirical cumulative distribution function (ECDF) of the covariates $\{\mathbf{X}_{i}\}_{i=1}^{n}$ as $F_{n}(\mathbf{x})=\sum_{i=1}^{n}\mathbb{I}(\mathbf{X}_{i}\leq\mathbf{x})/n$ and the weighted ECDF of the covariates for units assigned to treatment level $z\in\{0,1\}$, with weights $\mathbf{w}=(w_{1},\ldots,w_{n})^{\top}$, as $F_{\mathbf{w},z,n}(\mathbf{x})=\sum_{i=1}^{n}w_{i}\mathbb{I}(\mathbf{X}_{i}\leq\mathbf{x},Z_{i}=z)/n_{z}$. For notational brevity, we omit the subscript $i$ when referring to a generic unit, writing $(Y(z),Y,Z,\mathbf{X})$ in place of $(Y_{i}(z),Y_{i},Z_{i},\mathbf{X}_{i})$.
For a vector $\mathbf{v}$, we use $\|\mathbf{v}\|\equiv\|\mathbf{v}\|_{2}$ to denote its Euclidean ($\ell_{2}$) norm, and $\|\mathbf{v}\|_{1}$ and $\|\mathbf{v}\|_{\infty}$ to denote its $\ell_{1}$ and $\ell_{\infty}$ norms, respectively. For a sequence of random variables $\{T_{n}\}$ and a sequence of positive constants $\{a_{n}\}$, we write $T_{n}=O_{P}(a_{n})$ if $T_{n}/a_{n}$ is bounded in probability, and $T_{n}=o_{P}(a_{n})$ if $T_{n}/a_{n}\to 0$ in probability. Lastly, let $\mathcal{W}^{s,2}(\mathcal{V})$ be the (fractional) Sobolev space consisting of functions whose derivatives up to order $s$ have finite $L_{2}$ norm over a domain $\mathcal{V}$.
In this article, we focus on the estimation of the average treatment effect (ATE), defined by $\tau=\mathbb{E}\{Y(1)-Y(0)\}$. Extensions to other causal estimands in an IV setting are discussed in Section 5. In order to establish identification of the ATE, we make the following assumptions.
Assumption 1 (Causal Assumptions).
[…]
These assumptions are standard in causal inference; see Hernán and Robins (2020) for a detailed discussion. Under Assumption 1, it is well-known that the ATE can be identified by the following representations (Rosenbaum and Rubin, 1983; Hernán and Robins, 2020):
$\tau=\mathbb{E}\bigg[\bigg\{\frac{Z}{e(\mathbf{X})}-\frac{1-Z}{1-e(\mathbf{X})}\bigg\}Y\bigg]=\mathbb{E}\big\{\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X})\big\}\ ,$ (1)
where $e(\mathbf{X})$ is the propensity score and $\mu_{z}(\mathbf{X})\equiv\mathbb{E}(Y|Z=z,\mathbf{X})$ is the outcome regression.
2.2 Review: Weighting Estimators
The first representation of (1) serves as the key basis for weighting approaches, as it demonstrates that applying appropriate weights to treated and control units can eliminate confounding bias and allow for recovering causal effects. This principle—eliminating covariate imbalance via weighting—is the underlying idea behind these approaches. Formally, let $\mathbf{w}=(w_{1},\ldots,w_{n})^{\top}$ denote non-negative weights assigned to each unit. Given these weights, a weighting estimator of the ATE can be expressed as:
$\widehat{\tau}_{\mathbf{w}}=\frac{1}{n_{1}}\sum_{i=1}^{n}w_{i}Z_{i}Y_{i}-\frac{1}{n_{0}}\sum_{i=1}^{n}w_{i}(1-Z_{i})Y_{i}\ .$ (2)
Following Huling and Mak (2024), the error of $\widehat{\tau}_{\mathbf{w}}$ can be decomposed as:
$\widehat{\tau}_{\mathbf{w}}-\tau=$ $\int_{\mathcal{X}}\mu_{1}(\mathbf{x})d(F_{\mathbf{w},1,n}-F_{n})(\mathbf{x})-\int_{\mathcal{X}}\mu_{0}(\mathbf{x})d(F_{\mathbf{w},0,n}-F_{n})(\mathbf{x})$ (3)
$\quad-\int_{\mathcal{X}}\{\mu_{1}(\mathbf{x})-\mu_{0}(\mathbf{x})\}\,d(F-F_{n})(\mathbf{x})+\sum_{i=1}^{n}w_{i}\varepsilon_{i}\bigg(\frac{Z_{i}}{n_{1}}-\frac{1-Z_{i}}{n_{0}}\bigg),$ (4)
where $\varepsilon_{i}\equiv Y_{i}-\mu_{Z_{i}}(\mathbf{X}_{i})$, satisfying $\mathbb{E}(\varepsilon_{i}|Z_{i},\mathbf{X}_{i})=0$. Note that the terms in (4) converge to zero as the sample size grows. Consequently, any asymptotic bias of $\widehat{\tau}_{\mathbf{w}}$ as $n\rightarrow\infty$ arises from (3). In other words, the main source of potential bias in $\widehat{\tau}_{\mathbf{w}}$ is the discrepancy between $F_{\mathbf{w},z,n}$ and $F_{n}$, i.e., the weighted ECDFs and the unweighted ECDF. This underscores why the central aim in constructing weights is to minimize, ideally eliminate, imbalances in the covariate distributions of treated and control units relative to the marginal distribution.
We now provide a brief overview of how balancing weights can be constructed. The first approach is inverse probability weighting (IPW; Horvitz and Thompson, 1952; Hirano et al., 2003), which constructs weights based on the estimated propensity score $\widehat{e}$ and yields the following IPW estimator of the ATE:
$w_{i,\text{IPW}}=\frac{1}{n}\bigg\{\frac{n_{1}Z_{i}}{\widehat{e}(\mathbf{X}_{i})}+\frac{n_{0}(1-Z_{i})}{1-\widehat{e}(\mathbf{X}_{i})}\bigg\}\ ,\quad\widehat{\tau}_{\text{IPW}}=\frac{1}{n}\sum_{i=1}^{n}\left\{\frac{Z_{i}}{\widehat{e}(\mathbf{X}_{i})}-\frac{(1-Z_{i})}{1-\widehat{e}(\mathbf{X}_{i})}\right\}Y_{i}\ .$ (5)
The propensity score may be estimated using parametric methods (e.g., logistic regression) or nonparametric methods (e.g., machine learning). When the true propensity score is known, it can replace $\widehat{e}$ in the above expressions. A normalized version, known as the Hájek estimator (Hájek, 1971), rescales the weights so that they sum to one within treatment groups, i.e., $w_{i,\text{Hajek}}\propto w_{i,\text{IPW}}$ with $\sum_{i=1}^{n}\mathbb{I}(Z_{i}=z)w_{i,\text{Hajek}}=n_{z}$ for $z\in\{0,1\}$. The Hájek estimator often improves finite-sample stability compared to the IPW estimator.
The second approach is the covariate balancing propensity score (CBPS; Imai and Ratkovic, 2014), in which the propensity score is estimated by solving a set of moment equations that directly enforce covariate balance for a user-specified function $h(\cdot)$, often taken as the identity function. Specifically, the CBPS estimate $\widehat{e}_{B}$ is desiged to satisfy:
$0\;=\;\frac{1}{n}\sum_{i=1}^{n}\left\{\frac{Z_{i}}{\widehat{e}_{B}(\mathbf{X}_{i})}-\frac{1-Z_{i}}{1-\widehat{e}_{B}(\mathbf{X}_{i})}\right\}h(\mathbf{X}_{i})\ .$
By construction, the CBPS estimator $\widehat{e}_{B}$ is guaranteed to achieve covariate balance with respect to the specified basis function $h$, a property that the standard propensity score estimator $\widehat{e}$ in (5) may not generally possess. The resulting CBPS weighting estimator has the same form as (5), with $\widehat{e}_{B}$ replacing $\widehat{e}$.
The third approach is entropy balancing (Hainmueller, 2012; Chan et al., 2015), where the weights are obtained by solving a constrained optimization problem, say:
$\min_{\mathbf{w}}\sum_{i=1}^{n}h(w_{i})\quad\text{ subject to }\quad\bigg\{\begin{array}[]{l}\sum_{i=1}^{n}w_{i}c_{ri}(\mathbf{X}_{i})=m_{r}\quad\text{with }r\in\{1,\dots,R\}\\[-5.69046pt] \sum_{i=1}^{n}w_{i}\mathbb{I}(Z_{i}=z)=n_{z}\ ,\ z\in\{0,1\}\ ,\quad\mathbf{w}\geq 0\end{array}$
where $h(w_{i})$ is a distance metric between $w_{i}$ and reference weights (typically the uniform weights) and $c_{ri}(\mathbf{X}_{i})=m_{r}$ describes a set of $R$ balance constraints imposed on the covariate moments. By construction, entropy balancing guarantees that, after weighting, the specified moments of the covariates are exactly matched within each treatment group.
Lastly, a recent development in constructing weights is distributional balancing (Ben-Michael et al., 2021; Bruns-Smith et al., 2025; Chen et al., 2024; Clivio et al., 2024; Huling et al., 2024; Huling and Mak, 2024; Kallus, 2018, 2020; Kong et al., 2023; Wang et al., 2023), which aims to achieve balance across the full covariate distributions. In this approach, weights are obtained by minimizing a distance between two weighted empirical covariate distributions; see Section 3.2 for a formal mathematical characterization of these weights. Examples of such distances include maximum mean discrepancy (MMD; Gretton et al., 2012), energy distance (ED; Székely et al., 2004), the Kolmogorov-Smirnov statistic, $f$-divergences, and the Wasserstein distance. Unlike the aforementioned approaches, distributional balancing does not require specification of a propensity score model or a priori selection of moments to be balanced. Consequently, it is a nonparametric method that is robust to bias from model misspecification and can achieve covariate balance not only for selected moments but across the entire covariate distributions. As detailed in the remainder of this manuscript, our approach falls within the distributional balancing framework based on the characteristic function distance (CFD; Ansari et al., 2020; Heathcote, 1977; Chwialkowski et al., 2015).
3 Method: Unified Distributional Balancing
3.1 Characteristic Function Distance (CFD)
To introduce our approach, we first provide a brief review of the CFD. The CFD between two distributions $P$ and $Q$ with respect to a nonnegative density $\omega:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is defined as
$\texttt{CFD\hskip 1.05005pt}^{2}_{\omega}(P,Q)\;=\;\int_{\mathbb{R}^{d}}\big|\varphi_{P}(\mathbf{t})-\varphi_{Q}(\mathbf{t})\big|^{2}\,\omega(\mathbf{t})\,d\mathbf{t},$ (6)
where $\varphi_{P}(\mathbf{t})=\mathbb{E}_{\mathbf{V}\sim P}\{\exp({\sqrt{-1}\,\mathbf{t}^{\top}\mathbf{V}})\}$ is the characteristic function of $P$ (and similarly for $\varphi_{Q}$). The density function $\omega$ is assumed to satisfy certain properties, namely:
Assumption 2.
$\omega$ is nonnegative, finite, and integrable, with support equal to $\mathbb{R}^{d}$.
Under Assumption 2, the corresponding CFD defines a proper metric between distributions, as formalized in the following lemma, which is a direct consequence of Theorem 9 of Sriperumbudur et al. (2010):
Lemma 3.1.
Under Assumption 2, we have $\texttt{CFD\hskip 1.00006pt}_{\omega}^{2}(P,Q)=0$ if and only if $P=Q$.
Under Assumption 2, we may, without loss of generality, normalize $\omega$ so that it integrates to one, making it a probability density function on $\mathbb{R}^{d}$. While the finiteness and integrability requirements could be relaxed for certain cases (see Example 4 below), we retain them here for simplicity and to avoid unnecessary technical complications.
The CFD admits equivalent representations that highlight its connection to reproducing kernel Hilbert spaces (RKHS) and Fourier analysis. Given a density function $\omega$ satisfying Assumption 2, consider the translation-invariant kernel
$k_{\omega}(\mathbf{x},\mathbf{x}^{\prime})=\int_{\mathbb{R}^{d}}e^{\sqrt{-1}\,\mathbf{t}^{\top}(\mathbf{x}-\mathbf{x}^{\prime})}\,\omega(\mathbf{t})\,d\mathbf{t}\ ,$ (7)
which is well-defined and positive definite by Bochner’s theorem (Bochner, 2005), since $\omega$ acts as the spectral density of the kernel. In particular, $k_{\omega}$ is the inverse Fourier transform of $\omega$, and $\omega$ is the Fourier transform of $k_{\omega}$. Then, the CFD between $P$ and $Q$ can be expressed in kernel form as
$\texttt{CFD\hskip 1.05005pt}^{2}_{\omega}(P,Q)$ $=\mathbb{E}_{\mathbf{V},\mathbf{V}^{\prime}\sim P,\ \mathbf{W},\mathbf{W}^{\prime}\sim Q}\big\{k_{\omega}(\mathbf{V},\mathbf{V}^{\prime})+k_{\omega}(\mathbf{W},\mathbf{W}^{\prime})-2k_{\omega}(\mathbf{V},\mathbf{W})\big\},$ (8)
where $(\mathbf{V},\mathbf{V}^{\prime},\mathbf{W},\mathbf{W}^{\prime})$ are mutually independent with $\mathbf{V},\mathbf{V}^{\prime}\sim P$ and $\mathbf{W},\mathbf{W}^{\prime}\sim Q$. Furthermore, $\texttt{CFD\hskip 1.05005pt}_{\omega}^{2}$ admits a variational characterization as an MMD:
$\texttt{CFD\hskip 1.05005pt}^{2}_{\omega}(P,Q)=\left[\sup_{f\in\mathcal{H}_{\omega}:\,\|f\|_{\mathcal{H}_{\omega}}\leq 1}\big[\mathbb{E}_{\mathbf{V}\sim P,\mathbf{W}\sim Q}\{f(\mathbf{V})-f(\mathbf{W})\}\big]\right]^{2},$ (9)
where $\mathcal{H}_{\omega}$ is the RKHS associated with $k_{\omega}$, and $\|\cdot\|_{\mathcal{H}_{\omega}}$ is the corresponding RKHS norm. In other words, $\texttt{CFD\hskip 1.05005pt}_{\omega}^{2}(P,Q)$ is precisely the squared MMD between $P$ and $Q$ with respect to the kernel $k_{\omega}$. See Appendix LABEL:appendix-Equivalence_of_CFD_forms for details on the derivations of (8) and (9). In fact, the RKHS $\mathcal{H}_{\omega}$ is often characterized as a Sobolev space whose smoothness is determined by the choice of $\omega$; see Examples 1-4 below.
We now focus on the case where the two distributions of interest are the observed data, i.e., $\widehat{P}$ and $\widehat{Q}$ are the ECDFs of $\{\mathbf{V}_{i}\}_{i=1}^{n}\sim P$ and $\{\mathbf{W}_{j}\}_{i=1}^{m}\sim Q$, respectively. The goal is to evaluate the CFD between $\widehat{P}$ and $\widehat{Q}$. If the kernel function $k_{\omega}$ in (7) admits a closed-form representation, the CFD can be expressed in terms of the gram matrix as
[…]
However, obtaining a closed-form kernel may be infeasible when the density $\omega$ has a complex structure; see Examples 2 and 3. In such cases, the kernel (7) can be approximated via Monte Carlo. Specifically, for a sufficiently large $L$, let $\{\mathbf{T}_{\ell}\}_{\ell=1}^{L}$ be i.i.d. draws from a distribution with the probability density function proportional to $\omega$. Then, $k_{\omega}$ can be approximated by $\tilde{k}_{\omega}$, which serves as the empirical analogue of (7):
$\tilde{k}_{\omega}(\mathbf{x},\mathbf{x}^{\prime})$ $=\frac{1}{L}\sum_{\ell=1}^{L}e^{\sqrt{-1}\mathbf{T}_{\ell}^{\top}(\mathbf{x}-\mathbf{x}^{\prime})}.$ (11)
This approximation can be computed efficiently using the addition formula for trigonometric functions. Moreover, if $\omega$ is symmetric about zero, i.e., $\omega(\mathbf{t})=\omega(-\mathbf{t})$, then (11) reduces to an expression that involves only the real part; see Appendix LABEL:approximation_to_cos for details.
3.2 CFD-based Balancing Weights
As established in the preceding analysis, the primary source of potential bias in the weighting estimator $\widehat{\tau}_{\mathbf{w}}$ arises from the term in (3), which reflects the discrepancy between the weighted empirical distributions for the treated and control groups, $F_{\mathbf{w},1,n}$ and $F_{\mathbf{w},0,n}$, and the unweighted empirical distribution of the full sample, $F_{n}$. To reduce this bias, we seek weights $\mathbf{w}$ such that the weighted distributions $F_{\mathbf{w},z,n}$ become as similar as possible to the target distribution $F_{n}$ in terms of the CFD. This objective can be formalized as an optimization problem: minimize the sum of the CFDs between each treatment-group empirical distribution and the full-sample distribution. If this sum attains zero, the weighted empirical distributions are exactly equal to the target distribution from Lemma 3.1, i.e., perfectly balanced. Otherwise, the solution yields weights that render the distributions as similar as possible under the CFD metric.
Following this motivation, we define the CFD balancing weights as the solution to:
[…]
$\ \text{subject to}\quad\sum_{i=1}^{n}w_{i}Z_{i}=n_{1}\ ,\quad\sum_{i=1}^{n}w_{i}(1-Z_{i})=n_{0}\ ,\quad\mathbf{w}\geq 0\ .$ (12)
Here, $\lambda\geq 0$ is a regularization parameter introduced to prevent overfitting, which is required to decay at a certain rate with respect to $n$; see Assumption 7 for details. The constraints $\sum_{i=1}^{n}w_{i}Z_{i}=n_{1}$ and $\sum_{i=1}^{n}w_{i}(1-Z_{i})=n_{0}$ are included to normalize the weights within each treatment group. Their primary purpose is to ensure that the total “mass” or effective sample size of the weighted groups remains equal to that of the original, unweighted groups. This normalization is crucial for two reasons. First, it provides scale preservation, preventing the weights from arbitrarily inflating or deflating the influence of a group, which would distort the final estimate. Second, it leads to stable estimation, making the corresponding weighting estimators (e.g., $n_{1}^{-1}\sum w_{i}Y_{i}Z_{i}$) directly comparable in form and scale to their simple unweighted counterparts. This practice is analogous to the normalization of weights in the Hájek estimator reviewed in Section 2.2, which generally offers superior finite-sample properties relative to the IPW estimator.
Remark 1.
The formulation in (12) aims to find weights that make each treatment arm resemble the full sample but does not directly enforce similarity between the treatment and control groups. As a result, distributional discrepancies between the treated and control groups may remain, particularly in finite samples or when the two groups differ substantially in size or composition. To address this issue, one can adopt a “three-way” balancing objective that adds a direct discrepancy term between the two weighted groups. Following Chan et al. (2015) and Huling and Mak (2024), we consider
[…]
$\ \text{subject to}\quad\sum_{i=1}^{n}w_{i}Z_{i}=n_{1}\ ,\quad\sum_{i=1}^{n}w_{i}(1-Z_{i})=n_{0}\ ,\quad\mathbf{w}\geq 0\ .$ (13)
The additional term ${\texttt{CFD\hskip 1.05005pt}}_{\omega}^{2}(F_{\mathbf{w},1,n},F_{\mathbf{w},0,n})$ directly penalizes dissimilarity between the two groups’ empirical distributions under the CFD metric, aiming to achieve covariate balance in which the treated group, control group, and full sample are jointly aligned.
Now, we use (10) as a basis for solving the optimization tasks (12) and (13). Specifically, we take $\widehat{P}$ to be the weighted treated sample $F_{\mathbf{w},1,n}$ and $\widehat{Q}$ to be the unweighted full sample $F_{n}$. The weighted adaptation of the formula is:
[…]
$=\frac{1}{n_{1}^{2}}\mathbf{w}^{\top}D_{1}KD_{1}\mathbf{w}+\frac{1}{n^{2}}\mathbf{1}^{\top}K\mathbf{1}-\frac{2}{n_{1}n}\mathbf{w}^{\top}D_{1}K\mathbf{1}\ .$ (14)
In the first line, $\sum_{i,j}$ is shorthand $\sum_{i=1}^{n}\sum_{j=1}^{n}$. In the second line, $K\in\mathbb{R}^{n\times n}$ is the gram matrix of which $(i,j)$th entry is $k_{\omega}(\mathbf{X}_{i},\mathbf{X}_{j})$, $D_{1}=\text{diag}(Z_{1},\ldots,Z_{n})\in\mathbb{R}^{n\times n}$ is the diagonal matrix indicating the treated units, and $\mathbf{1}=(1,\ldots,1)^{\top}\in\mathbb{R}^{n}$ is a vector of ones. Similarly, taking $(\widehat{P},\widehat{Q})=(F_{\mathbf{w},0,n},F_{n})$ and $(\widehat{P},\widehat{Q})=(F_{\mathbf{w},1,n},F_{\mathbf{w},0,n})$ yields
$\texttt{CFD\hskip 1.05005pt}^{2}_{\omega}(F_{\mathbf{w},0,n},F_{n})=\frac{1}{n_{0}^{2}}\mathbf{w}^{\top}D_{0}KD_{0}\mathbf{w}+\frac{1}{n^{2}}\mathbf{1}^{\top}K\mathbf{1}-\frac{2}{n_{0}n}\mathbf{w}^{\top}D_{0}K\mathbf{1},$ (15)
$\texttt{CFD\hskip 1.05005pt}^{2}_{\omega}(F_{\mathbf{w},1,n},F_{\mathbf{w},0,n})=\frac{1}{n_{1}^{2}}\mathbf{w}^{\top}D_{1}KD_{1}\mathbf{w}-\frac{2}{n_{1}n_{0}}\mathbf{w}^{\top}D_{1}KD_{0}\mathbf{w}+\frac{1}{n_{0}^{2}}\mathbf{w}^{\top}D_{0}KD_{0}\mathbf{w},$ (16)
where $D_{0}=\text{diag}(1-Z_{1},\ldots,1-Z_{n})\in\mathbb{R}^{n\times n}$ is the diagonal matrix indicating the control units. Substituting (14)-(16) into (13), we obtain
$\mathbf{w}_{\texttt{CFD\hskip 0.73503pt}}\in\operatorname*{arg\,min}_{\mathbf{w}\in\mathbb{R}^{n}}\big(\mathbf{w}^{\top}\mathcal{Q}\mathbf{w}+\mathbf{q}^{\top}\mathbf{w}\big)\ \ \text{s.t.}\ \ \sum_{i=1}^{n}w_{i}Z_{i}=n_{1},\ \sum_{i=1}^{n}w_{i}(1-Z_{i})=n_{0},\ \mathbf{w}\geq 0,$ (17)
$\mathcal{Q}=\frac{D_{1}KD_{1}}{n_{1}^{2}}+\frac{D_{0}KD_{0}}{n_{0}^{2}}-\frac{D_{1}KD_{0}}{n_{1}n_{0}}+\lambda^{2}I_{n}\ ,\quad\mathbf{q}=-\left(\frac{D_{1}K\mathbf{1}}{n_{1}n}+\frac{D_{0}K\mathbf{1}}{n_{0}n}\right)\ .$
A similar formulation can be derived for the case of (12). Equation (17) is a standard constrained quadratic programming (QP) task with a unique global solution, as the objective function is strictly convex for $\lambda>0$ (Boyd and Vandenberghe, 2004). This problem can be solved using general-purpose QP solvers; for instance, we used an R library optiSolve (Wellmann, 2021) in the simulation and data analysis; see Appendix LABEL:sec-QP on details of how this is implemented. Of note, when the closed-form representation of the CFD is not available, the approximated kernel in (11) can be used in (14)-(17).
[…]
$\widehat{\tau}_{\texttt{CFD\hskip 0.73503pt}}=\frac{1}{n_{1}}\sum_{i=1}^{n}w_{i,\texttt{CFD\hskip 0.73503pt}}Z_{i}Y_{i}-\frac{1}{n_{0}}\sum_{i=1}^{n}w_{i,\texttt{CFD\hskip 0.73503pt}}(1-Z_{i})Y_{i}\ .$ (18)
3.3 Unification: Examples of Our Framework
Many distributional distances can be expressed in the CFD form of (6) with a suitable choice of the density function $\omega$. Consequently, our approach provides a unified framework for existing distributional balancing methods. Several examples are discussed below.
Example 1 (Gaussian Density).
A prominent example is the MMD with a Gaussian kernel, which arises when $\omega(\mathbf{t})$ is a multivariate Gaussian density, specifically $\omega_{\text{G}}(\mathbf{t})\propto\exp(-\gamma^{2}\|\mathbf{t}\|_{2}^{2}/4)$ for a bandwidth parameter $\gamma>0$. The inverse Fourier transform of this Gaussian density yields the Gaussian kernel, denoted by $k_{\text{G}}(\mathbf{x},\mathbf{x}^{\prime})=\exp(-\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}/\gamma^{2})$. The corresponding RKHS $\mathcal{H}_{\text{G}}$ is a collection of functions that are exceptionally smooth; specifically, $\mathcal{H}_{\text{G}}$ is contained in the intersection of infinitely many Sobolev spaces: $\mathcal{H}_{\text{G}}\subset\mathcal{W}^{s,2}(\mathcal{X})$ for any $s>0$ (Steinwart and Christmann, 2008, Corollary 4.36), meaning that every function in $\mathcal{H}_{\text{G}}$ is infinitely differentiable.
Example 2 (Separable Density).
Another common choice for the density function is a separable one, which can be written as a product of one-dimensional densities. The $\ell_{1}$-Laplacian kernel provides a key instance of this case. It corresponds to selecting $\omega_{\text{L}}(\mathbf{t})\propto\prod_{j=1}^{d}(1+\gamma^{2}t_{j}^{2})^{-1}$, which is a product of one-dimensional Cauchy densities. The resulting kernel is the $d$-dimensional $\ell_{1}$-Laplacian kernel, $k_{\text{L}}(\mathbf{x},\mathbf{x}^{\prime})=\exp(-\|\mathbf{x}-\mathbf{x}^{\prime}\|_{1}/\gamma)=\prod_{j=1}^{d}\exp(-|x_{j}-x_{j}^{\prime}|/\gamma)$, which itself is a product of one-dimensional Laplacian kernels.
The one-dimensional kernel $\exp(-|x_{j}-x_{j}^{\prime}|/\gamma)$ has the one-dimensional Sobolev space $\mathcal{W}^{1,2}(\mathcal{X}_{j})$ as its native space (Wendland, 2004) where $\mathcal{X}_{j}$ is the support of $X_{j}$. Since $k_{\text{L}}$ is a tensor product of these one-dimensional kernels, its native space is the tensor product Sobolev space, i.e., $\mathcal{H}_{\text{L}}=\bigotimes_{j=1}^{d}\mathcal{W}^{1,2}(\mathcal{X}_{j})$ (Berlinet and Thomas-Agnan, 2011, Theorem 13).
The above Cauchy density can be replaced with generic one-dimensional densities. For instance, we can select $\omega_{\text{S},j}(t_{j})$ to be a scaled density of the Student’s t-distribution with $(2s_{j}-1)$ degrees of freedom, which corresponds to $\omega_{\text{S},j}(t_{j})\propto(1+\gamma^{2}t_{j}^{2})^{-s_{j}}.$ For a certain $s_{j}$, the corresponding one-dimensional kernel $k_{\text{S},j}$ has a known closed form as a Matérn kernel (Rasmussen and Williams, 2006). The native space for the one-dimensional kernel $k_{\text{S},j}$ is the Sobolev space $\mathcal{W}^{s_{j},2}(\mathcal{X}_{j})$, and consequently, the native space for the full kernel $k_{\text{S}}$ is $\mathcal{H}_{\text{S}}=\bigotimes_{j=1}^{d}\mathcal{W}^{s_{j},2}(\mathcal{X}_{j})$. Notably, the Laplacian kernel $k_{\text{L}}$ and the Gaussian kernel $k_{\text{G}}$ in Example 1 correspond to the special cases $s_{j}=1$ and $s_{j}=\infty$, respectively, for all $j$. Therefore, specifying the Laplacian and Gaussian kernels can be seen as two extreme cases, while the general case with $s_{j}\in(1,\infty)$ represents an intermediate one between them.
In fact, $\omega_{\text{S},j}$ does not need to be exactly proportional to the t-distribution density. Specifically, suppose there exist positive constants $c_{1}$ and $c_{2}$ such that $c_{1}(1+\gamma^{2}t_{j}^{2})^{-s_{j}}\leq\omega_{\text{S},j}(t_{j})\leq c_{2}(1+\gamma^{2}t_{j}^{2})^{-s_{j}}$. This relaxation allows the density $\omega_{\text{S},j}(t_{j})$ to be complex, provided that its tail behavior resembles that of a t-distribution. In such cases, obtaining a closed-form expression for the coordinate-wise and full kernels, $k_{\text{S},j}$ and $k_{\text{S}}$, may be challenging. Nevertheless, these kernels can be effectively approximated using Monte Carlo integration via (11). Importantly, despite this relaxation, the native space associated with the full kernel $k_{\text{S}}$ remains $\mathcal{H}_{\text{S}}=\otimes_{j=1}^{d}\mathcal{W}^{s_{j},2}(\mathcal{X}_{j})$ (Wendland, 2004, Corollary 10.13).
Example 3 (Isotropic Density).
Our framework also includes isotropic (i.e., rotationally invariant) densities. For example, consider the density $\omega_{\text{M}}(\mathbf{t})\propto(1+\gamma^{2}\|\mathbf{t}\|_{2}^{2})^{-s}$, where $s$ is a smoothness parameter with $s>d/2$. Unlike the previous example, this density cannot be decomposed into a product of one-dimensional functions. Consequently, the corresponding kernel $k_{\text{M}}$ is a $d$-dimensional Matérn kernel rather than the product of one-dimensional Matérn kernel; see Appendix LABEL:Spectral_Densities_for_Different_Kernels for the closed-form representation of it. The resulting native space $\mathcal{H}_{\text{M}}$ is a $d$-dimensional Sobolev space $\mathcal{W}^{s,2}(\mathcal{X})$ (Wendland, 2004, Chapter 10), which differs from $\mathcal{H}_{\text{S}}$ in the previous example. As in the previous example, this condition can be relaxed: specifically, if there exist positive constants $c_{1}$ and $c_{2}$ such that $c_{1}(1+\gamma^{2}\|\mathbf{t}\|_{2}^{2})^{-s}\leq\omega_{\text{M}}(\mathbf{t})\leq c_{2}(1+\gamma^{2}\|\mathbf{t}\|_{2}^{2})^{-s}$, then the resulting native space remains $\mathcal{H}_{\text{M}}=\mathcal{W}^{s,2}(\mathcal{X})$ (Wendland, 2004, Corollary 10.13).
Example 4 (Energy Distance).
While primarily focused on proper densities $\omega$ satisfying Assumption 2, our framework can also accommodate certain “improper densities,” with the ED serving as a key example. For the ED, the associated density is $\omega_{\text{ED}}(\mathbf{t})\propto\|\mathbf{t}\|_{2}^{-(d+1)}$, which does not satisfy Assumption 2. Consequently, the corresponding kernel cannot be obtained via the standard inverse Fourier transform in (7). Nevertheless, the inverse Fourier transform exists in a distributional sense, yielding the energy distance kernel $k_{\text{ED}}(\mathbf{x},\mathbf{x}^{\prime})=-\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}$ (Székely and Rizzo, 2013). Note that this kernel is not positive definite, as Bochner’s theorem does not apply. However, following Mak and Joseph (2018), the associated native space $\mathcal{H}_{\text{ED}}$ can still be characterized and is closely related to Sobolev spaces. Specifically, the Sobolev space $\mathcal{W}^{(d+1)/2,2}$ is embedded in $\mathcal{H}_{\text{ED}}$, namely $\mathcal{W}^{(d+1)/2,2}\subset\mathcal{H}_{\text{ED}}$. Moreover, when the dimension $d$ is odd, the spaces are exactly equal: $\mathcal{W}^{(d+1)/2,2}=\mathcal{H}_{\text{ED}}$. This provides a unified mathematical representation of the ED as a type of CFD, since the formulations in (6), (8), and (9) remain valid for the ED, even though it is not a standard CFD.
Remark 2 (Exception: Wasserstein distance).
Another popular discrepancy metric is the Wasserstein $p$-distance ($W_{p}$). While the discussion can be extended to general $p$, we focus on $p=1$ as a motivating example. Note that the $W_{1}$ distance is represented as $W_{1}(P,Q)=\sup_{f\in\mathcal{F}}[\mathbb{E}_{\mathbf{V}\sim P,\mathbf{W}\sim Q}\{f(\mathbf{V})-f(\mathbf{W})\}]$, where $\mathcal{F}$ is the class of 1-Lipschitz functions. This representation shares similarity as the CFD in (9), except that the function classes defining the distances differ. However, Modeste and Dombry (2024) showed that the Wasserstein distance cannot be represented in the CFD form of (6). Consequently, the Wasserstein distance is a notable exception that our method does not unify.
Nonetheless, this exception does not diminish the practical advantages of our approach. The native space for the $W_{1}$ distance (i.e., 1-Lipschitz functions) is significantly more complex than the Sobolev spaces discussed above, as demonstrated by a formal comparison of their covering numbers in Appendix LABEL:Comparison_between_different_Function_Classes. A direct consequence of this complexity is that the $W_{1}$ distance suffers from the curse of dimensionality; specifically, it is well-known that $W_{1}(F_{n},F)=O_{P}(n^{-1/d})$ for $d\geq 3$ (van Handel, 2014, page 137), leading to substantially slower convergence than the $\sqrt{n}$-rate. Therefore, even when the corresponding balancing weights are well-calibrated and the outcome regressions ($\mu_{z}$) are sufficiently smooth, using the $W_{1}$ distance may yield a weighting estimator that fails to achieve $\sqrt{n}$-consistency. In contrast, our unified CFD-based approach produces a $\sqrt{n}$-consistent weighting estimator as long as $\mu_{z}$ exhibits an appropriate degree of smoothness relative to the specified density $\omega$. More specifically, if $\mu_{z}$ belongs to the native space corresponding to $\omega$, $\sqrt{n}$-consistency can be established; see Theorem 4.1 for details.
4 Theory
4.1 $\sqrt{n}$-Consistency of the Weighting Estimator
We now present the statistical properties of the weighting estimator, where the weights are obtained from the CFD balancing problem in (13). We note that the statistical properties of the weighting estimator based on (12) can be established analogously. We begin by specifying the required assumptions.
Assumption 3 (Compact Covariate Support).
The covariate support $\mathcal{X}$ is compact.
Assumption 3 ensures that uniform boundedness and convergence of integrals involving kernels, distance metrics, and characteristic functions.
Assumption 4 (Smoothness of the Outcome Regression).
$\mu_{z}(\cdot)\in\mathcal{H}_{\omega}$ for $z\in\{0,1\}$.
Assumption 4 requires the true outcome regressions $\mu_{z}$ to lie within the native space $\mathcal{H}_{\omega}$. This is a critical smoothness assumption because properties of $\mathcal{H}_{\omega}$ are entirely determined by the choice of $\omega$. This directly connects to one of our contributions, which highlights that the tail behavior of $\omega$ dictates the smoothness of the functions contained in its native space. For instance, a rapidly decaying $\omega$ (e.g., Gaussian) induces a native space consisting of exceptionally smooth functions, thereby imposing a very strong condition on $\mu_{z}$. In contrast, a heavy-tailed $\omega$ (e.g., Laplacian) generates a native space that contains less smooth functions, corresponding to a weaker, less restrictive assumption. Thus, the choice of $\omega$ implicitly specifies the smoothness required of the true outcome model.
Assumption 5 (Boundeness of the Outcome).
For $z\in\{0,1\}$, $\text{Var}\{\mu_{z}(\mathbf{X})\}<\infty$ and $\sigma_{z}^{2}(\mathbf{X})\equiv\text{Var}(Y\,|\,Z=z,\mathbf{X})$ is bounded over $\mathbf{X}\in\mathcal{X}$.
Assumption 5 imposes conditions on the outcome distribution; these conditions are satisfied when $Y$ is uniformly bounded.
Assumption 6 (Stable Weights).
[…]
Assumption 6 states that the estimated balancing weights are not “too large”, specifically that no single weight grows faster than an $n^{1/3}$-rate. Several weighting-based methods for covariate balancing rely on this assumption (Wong and Chan, 2018; Athey et al., 2018; Huling and Mak, 2024). It prevents a pathological scenario where one or a few data points receive enormous weight (e.g., weights on the order of $n$), which would make the estimator highly unstable by depending on just a few observations.
Assumption 7 (Regularization Rate).
The regularization parameter in (13) is chosen with a rate of $\lambda=o(n^{-1})$.
Assumption 7 guarantees that the penalty term in (13), $\lambda^{2}\|\mathbf{w}\|_{2}^{2}$, becomes asymptotically negligible relative to the CFD terms.
Under these assumptions, the CFD-based weighting estimator in (18) is $\sqrt{n}$-consistent:
Theorem 4.1.
[…]
Leveraging the result in Theorem 4.1, we can construct confidence intervals for inference; see Section 4.2 for details. We emphasize that Assumption 2 is a sufficient condition used to simplify the presentation; in some cases, $\sqrt{n}$-consistency can still hold even when it is violated.
Remark 3.
Although the density $\omega$ can be chosen to satisfy Assumption 2, it is also of interest to consider choices that do not. In fact, Theorem 4.1 may still hold even when Assumption 2 is violated. For example, consider the ED density from Example 4, $\omega_{\text{ED}}(\mathbf{t})=\|\mathbf{t}\|_{2}^{-(d+1)}$. Although this choice clearly violates Assumption 2, the corresponding weighting estimator remains $\sqrt{n}$-consistent; see Huling and Mak (2024) for a formal proof of this result. Additional intuition is provided in Remark LABEL:remark:_relaxation_of_omega_for_ED in the Appendix.
Remark 4.
A consequence of semiparametric efficiency theory (Bickel et al., 1998) is that any regular and asymptotically linear (RAL) estimator of the ATE in the nonparametric model, denoted by $\widehat{\tau}_{\text{RAL}}$, admits the following unique asymptotic representation:
$\sqrt{n}(\widehat{\tau}_{\text{RAL}}-\tau)$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\underbrace{\bigg[\frac{(2Z_{i}-1)\{Y_{i}-\mu_{Z_{i}}(\mathbf{X}_{i})\}}{Z_{i}e(\mathbf{X}_{i})+(1-Z_{i})\{1-e(\mathbf{X}_{i})\}}+\mu_{1}(\mathbf{X}_{i})-\mu_{0}(\mathbf{X}_{i})-\tau\bigg]}_{\texttt{IF}(Y_{i},Z_{i},\mathbf{X}_{i})}+o_{P}(1),$ (19)
[…]
4.2 Inference via Subsampling
Bootstrap (Efron and Tibshirani, 1994) is a widely used method and is often regarded as a readily accessible tool for conducting inference in practice. Despite its popularity, a theoretically important but less frequently discussed point is that its validity generally relies on the estimator being regular, i.e., it can fail for non-regular estimators (Shao, 1994; Abadie and Imbens, 2008; Andrews, 2000; Chakraborty et al., 2010). As discussed in Remark 4, our estimator may not be regular, implying that the bootstrap procedure may not yield valid statistical inference. Indeed, we observe in our simulations that the standard bootstrap can fail to provide valid confidence intervals; see Section 6.
As a valid alternative, we consider subsampling. Specifically, Politis and Romano (1994) showed that subsampling remains valid for a wide class of estimators, including non-regular ones, as long as the estimator’s rate of convergence is known. Since we established the $\sqrt{n}$-consistency of the proposed weighting estimator in Theorem 4.1, subsampling provides a valid method for constructing confidence intervals. Nonetheless, subsampling often produces conservative inference, even for large sample sizes, a phenomenon commonly observed in practice. For theoretical explanations of this conservativeness, see Andrews and Guggenberger (2010). A practical challenge in implementing subsampling is the choice of the subsample size. To address this, we propose a data-driven method for selecting the subsample size following the volatility-based procedure of Politis et al. (1999). The full details of this selection algorithm are provided in Appendix LABEL:appendix-subsampling.
5 Extension to an Instrumental Variable Setting
The proposed method relies on the no unmeasured confounding assumption, namely Assumption 1(ii), but it can be extended to account for unmeasured confounding. Specifically, we consider the IV framework of Angrist et al. (1996). Let $\mathbf{X}\in\mathcal{X}$ denote the vector of observed pre-treatment covariates, $Z\in\{0,1\}$ the binary instrument, $A\in\{0,1\}$ the binary indicator of treatment receipt, where $A=1$ indicates that a unit receives the treatment and $A=0$ otherwise, and $Y\in\mathbb{R}$ the outcome variable. Let $A(z)$ denote the potential treatment status under $Z=z$, $Y(z,a)$ the potential outcome under $(Z=z,A=a)$, and $Y(z)=Y(z,A(z))$ the potential outcome under $Z=z$ with $A$ set to the natural treatment receipt. We impose the following IV assumptions, as detailed in Angrist et al. (1996) and Levis et al. (2024):
Assumption 8 (Instrumental Variable).
[…]
In this context, a primary causal estimand of interest is the local average treatment effect (LATE), defined as $\tau_{\text{LATE}}=\mathbb{E}\{Y(a=1)-Y(a=0)\mid A(z=1)>A(z=0)\}$, assuming $\mathrm{Pr}\{A(z=1)>A(z=0)\}>0$ to ensure that the LATE is well-defined. Under Assumptions 1(iii) and 8, the LATE is identified as:
$\tau_{\text{LATE}}=\frac{\mathbb{E}\{\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X})\}}{\mathbb{E}\{\nu_{1}(\mathbf{X})-\nu_{0}(\mathbf{X})\}}\ ,\ \mu_{z}(\mathbf{X})=\mathbb{E}(Y\,|\,Z=z,\mathbf{X})\ ,\ \nu_{z}(\mathbf{X})=\mathbb{E}(A\,|\,Z=z,\mathbf{X})\ .$ (20)
[…]
$\widehat{\tau}_{\text{LATE},\texttt{CFD\hskip 0.73503pt}}=\frac{\frac{1}{n_{1}}\sum_{i=1}^{n}Z_{i}w_{i,\texttt{CFD\hskip 0.73503pt}}Y_{i}-\frac{1}{n_{0}}\sum_{i=1}^{n}(1-Z_{i})w_{i,\texttt{CFD\hskip 0.73503pt}}Y_{i}}{\frac{1}{n_{1}}\sum_{i=1}^{n}Z_{i}w_{i,\texttt{CFD\hskip 0.73503pt}}A_{i}-\frac{1}{n_{0}}\sum_{i=1}^{n}(1-Z_{i})w_{i,\texttt{CFD\hskip 0.73503pt}}A_{i}}.$ (21)
[…]
Assumption 9 (Smoothness of Nuisance Functions).
For $z\in\{0,1\}$, both the outcome regression $\mu_{z}(\cdot)$ and the treatment regression $\nu_{z}(\cdot)$ belong to $\mathcal{H}_{\omega}$.
Theorem 5.1.
[…]
6 Simulation
We conduct a simulation study to evaluate the finite-sample performance of the proposed CFD-based weighting estimators. To illustrate our method’s performance beyond the standard no unmeasured confounding scenario, we consider the following data-generating process under the IV setting outlined in Section 5. First, we fix the sample size to $n\in\{100,200,400,800,1600\}$. For each subject $i\in\{1,\ldots,n\}$, we generate a $10$-dimensional pre-treatment covariate $\mathbf{X}_{i}=(X_{i1},\ldots,X_{i10})^{\top}$, whose components are independent standard normal random variables, i.e., $\mathbf{X}_{i}\sim N(\mathbf{0},I_{10})$. We also generate a variable $U_{i}\sim N(0,1)$, independently from $\mathbf{X}_{i}$. We treat $\mathbf{X}_{i}$ and $U_{i}$ as measured and unmeasured confounders, respectively.
The binary instrument $Z_{i}$ is generated from $Z_{i}\sim\text{Ber}(e(\mathbf{X}_{i}))$, where the propensity score $e$ is specified according to one of two models: (linear) $e_{\text{Linear}}(\mathbf{X}_{i})=\text{expit}\big(0.15\sum_{j=1}^{10}X_{ij}\big)$; or (nonlinear) $e_{\text{Nonlinear}}(\mathbf{X}_{i})=\text{expit}\big(0.25\sum_{j=1}^{5}|X_{ij}|+0.25\sum_{j=6}^{10}X_{ij}-2.25\big)$, where $\text{expit}(v)=\exp(v)/\{1+\exp(v)\}$. The potential treatment receipts under each instrument level, $A_{i}(0)$ and $A_{i}(1)$, are generated as $A_{i}(0)=\mathbbm{1}(L_{1i}-0.5|U_{i}|\geq 0)$ and $A_{i}(1)=\mathbbm{1}(L_{1i}+0.5|U_{i}|\geq 0)$ where $L_{1i}=-0.01-0.5\sum_{j=1}^{5}X_{ij}+0.5\sum_{j=6}^{10}X_{ij}$. This specification ensures that the monotonicity assumption holds, i.e., $A_{i}(1)\geq A_{i}(0)$ for all $i$. The potential outcomes are generated as $Y_{i}(a)=a(1+L_{2i})-0.5L_{2i}+\varepsilon_{ai}$ for $a\in\{0,1\}$ where $L_{2i}=\sum_{j=1}^{5}X_{ij}-0.5\sum_{j=6}^{10}X_{ij}-U_{i}$ and noise terms $\varepsilon_{ai}$ are independently drawn from $N(0,1)$. The observed treatment and outcome are given by $A_{i}=A_{i}(Z_{i})$ and $Y_{i}=Y_{i}(A_{i})$. In the analysis, we use $(Y_{i},A_{i},Z_{i},\mathbf{X}_{i})$, while $U_{i}$ is omitted as it serves as an unmeasured confounder.
We estimate the LATE using our proposed CFD-based weighting estimators with five different specifications of the density function $\omega$: (i) Gaussian: $\omega_{\text{G}}$ from Example 1, which yields the Gaussian kernel; (ii) $\ell_{1}$-Laplacian: $\omega_{\text{L}}$ from Example 2, producing the $\ell_{1}$-Laplacian kernel; (iii) t5: $\omega_{\text{S}}$ from Example 2, where $\omega_{\text{S},j}$ is proportional to a t-distribution with 5 degrees of freedom; (iv) Matérn: $\omega_{\text{M}}$ from Example 3 with $s=7.5$, producing the Matérn kernel; and (v) energy: $\omega_{\text{ED}}$ from Example 4, corresponding to the energy distance. Except for (iii), the CFD is evaluated using the closed-form representation of the kernel $k_{\omega}$, whereas for (iii), it is approximated using (11) with $L=10^{4}$ Monte Carlo repetitions. For (i)-(iv), the density requires selecting bandwidth hyperparameters $\gamma$ (see Examples 1-3), which we choose using the median heuristic (Garreau et al., 2018). For all proposed estimators, we use the three-way balancing in Remark 1, and the regularization parameter in (13) is set to $\lambda=n^{-2}$, which satisfies Assumption 7. In addition, we construct 95% confidence intervals using both subsampling and bootstrap methods for inference. As competing methods, we also consider weighting estimators based on weights derived from the parametric IPW and CBPS approaches described in Section 2.2. For the IPW estimator, $e(\cdot)$ is estimated using logistic regression of $Z$ on $\mathbf{X}$ with main-effect terms only; for CBPS, we use the implementation in the WeightIt R package (Greifer, 2025), which estimates propensity score weights under the same logistic specification while targeting balance of covariate means. Further details on the median heuristic, the construction of confidence intervals, and the implementation of the IPW and CBPS estimators are provided in Appendices LABEL:appendix-median-heuristic-LABEL:implementation_of_the_IPW_and_CBPS_estimators. We repeat the simulation 500 times, computing the bias and empirical standard error (ESE) for all seven methods. For inference, we report the empirical coverage rates of the 95% subsampling and bootstrap confidence intervals for our weighting estimators (i)-(v), and those of bootstrap confidence intervals for the IPW and CBPS estimators.
Table 1 presents a summary of the simulation results. We begin by examining the bias and ESE of the competing estimators. In both the linear and non-linear propensity score model scenarios, the bias and ESE of our weighting estimators all decrease steadily toward zero as the sample size $n$ increases, and the decrease appears to follow an $n^{-1/2}$-rate as demonstrated in Appendix LABEL:sec:validation_of_root_n_consistency. This result is consistent with Theorem 4.1. In contrast, the IPW and CBPS methods show a critical vulnerability to model misspecification. While these estimators appear consistent under the linear propensity score setting, they fail in the non-linear propensity score scenario, as their bias remains large and persistent regardless of sample size. This failure is expected, given the severe misspecification of their underlying parametric propensity score models.
Next, we compare the subsampling and bootstrap inference procedures. Across all scenarios and for all five weighting estimators, the subsampling confidence intervals attain the nominal 95% coverage level. This confirms that subsampling provides valid, albeit slightly conservative, inference for our estimators. In contrast, the bootstrap confidence intervals perform poorly, particularly for (ii) the $\ell_{1}$-Laplacian and (v) energy distance cases. This finding empirically supports our earlier concern regarding the failure of the standard bootstrap for the weighting estimator, as discussed in Section 4.2. While the subsampling intervals are consistently wider than the bootstrap intervals, this moderate increase in width is reasonable and reflects the method’s conservative nature, which guards against undercoverage. In comparison, the bootstrap intervals may fail dramatically, providing misleadingly narrow intervals with substantial coverage deficits. Consequently, we recommend subsampling as the preferred method for inference. The results clearly demonstrate the statistical properties and robustness to model misspecification of the proposed CFD-based weighting estimators, while also confirming subsampling as a reliable inferential procedure.
Method NN Data-generating process
Linear Nonlinear
Bias ESE Coverage Length Bias ESE Coverage Length
$S$ Boot $S$ Boot $S$ Boot $S$ Boot
Distributional Balancing Example 1 Gaussian 100 0.9 62.6 97.8 94.0 4.5 2.6 -5.7 67.6 95.8 92.0 5.0 2.8
200 0.9 43.4 98.0 95.4 2.4 1.6 1.2 44.3 98.0 94.0 2.5 1.6
400 1.2 30.5 98.0 96.4 1.5 1.1 1.2 30.2 98.2 95.8 1.5 1.1
800 -1.3 20.4 98.6 98.2 1.0 0.8 -0.2 22.7 97.8 96.2 1.0 0.8
1600 -0.4 16.0 96.4 96.6 0.6 0.6 -0.5 15.2 98.6 99.2 0.7 0.6
Example 2 $\ell_{1}$-Laplacian 100 1.9 65.5 97.8 90.8 5.0 2.5 -13.1 72.5 94.8 88.0 5.7 2.8
200 0.4 42.2 97.2 91.8 2.6 1.6 -6.2 46.2 96.8 89.8 2.8 1.6
400 0.9 28.1 98.0 91.4 1.5 1.0 -1.5 30.5 95.8 91.4 1.6 1.0
800 -0.7 19.7 97.8 89.4 1.0 0.7 -1.4 20.0 97.2 90.2 1.0 0.7
1600 -0.4 13.2 98.2 90.4 0.7 0.4 -1.5 12.8 99.0 91.8 0.7 0.5
Example 2 $t_{5}$ 100 0.2 70.9 99.6 96.4 6.2 3.5 -3.6 76.3 98.4 95.4 6.5 3.7
200 1.1 49.3 99.2 96.6 3.2 2.0 4.0 52.1 99.2 96.2 3.3 2.0
400 0.5 33.1 99.6 97.6 1.8 1.3 3.6 33.6 99.4 97.8 1.9 1.3
800 -0.6 20.7 99.4 97.8 1.1 0.9 2.5 22.9 99.4 97.0 1.1 0.9
1600 -0.2 13.9 98.8 96.6 0.7 0.6 0.5 13.7 99.8 98.0 0.7 0.6
Example 3 Matérn with $s=7.5$ 100 1.2 61.5 98.0 93.4 4.5 2.6 -6.5 66.7 95.8 91.2 5.1 2.8
200 0.7 41.2 97.8 94.4 2.4 1.5 0.4 42.5 97.2 93.4 2.5 1.6
400 1.4 28.3 98.4 94.2 1.5 1.0 0.8 28.2 98.6 93.8 1.5 1.0
800 -1.1 18.2 98.8 96.4 0.9 0.7 -0.4 20.1 98.4 94.2 1.0 0.7
1600 -0.8 13.5 97.6 94.2 0.6 0.5 -0.7 13.2 99.4 95.4 0.6 0.5
Example 4 Energy 100 2.1 62.4 97.8 92.8 4.4 2.4 -12.0 68.6 94.6 88.8 4.9 2.6
200 0.9 39.2 97.6 91.4 2.4 1.5 -4.7 42.4 96.0 90.2 2.5 1.5
400 1.0 25.5 98.8 90.4 1.4 0.9 -2.1 27.5 97.6 90.0 1.5 0.9
800 -0.8 17.2 98.0 89.2 0.9 0.6 -1.6 18.3 98.0 88.4 1.0 0.6
1600 -0.7 11.8 98.0 89.2 0.6 0.4 -1.4 11.8 98.8 90.2 0.6 0.4
IPW 100 8.5 103.3 - 99.0 - 7.6 -35.4 80.2 - 95.0 - 5.8
200 0.2 52.4 - 96.0 - 2.5 -33.4 53.0 - 91.0 - 2.2
400 0.0 34.2 - 96.0 - 1.5 -32.3 38.0 - 83.0 - 1.4
800 -0.3 26.5 - 92.8 - 1.0 -33.2 25.7 - 71.6 - 1.0
1600 -0.5 17.5 - 94.6 - 0.7 -33.1 17.0 - 52.8 - 0.7
CBPS 100 5.4 74.4 - 96.6 - 3.7 -31.7 79.8 - 93.2 - 3.8
200 -0.3 50.6 - 95.6 - 2.1 -31.0 53.3 - 90.8 - 2.2
400 -0.1 33.4 - 96.2 - 1.4 -31.2 38.1 - 84.0 - 1.4
800 -0.2 26.2 - 92.6 - 1.0 -32.6 25.7 - 73.6 - 1.0
1600 -0.5 17.4 - 94.6 - 0.7 -32.8 17.0 - 54.4 - 0.7
Table 1: Summary of the simulation studies. Bias and ESE are scaled by a factor of 100.
7 Application
We apply the proposed method to a real-world setting, following the analysis of Chernozhukov et al. (2018), and focus on the effect of 401(k) participation on financial asset accumulation. We use data from the DoubleML R package (Bach et al., 2022), which includes $n=9,915$ individuals, of whom $3,682$ are eligible for a 401(k) plan. The primary variables are net financial assets ($Y$), 401(k) participation ($A$), and 401(k) eligibility ($Z$). To adjust for confounding, we include nine pre-treatment covariates ($\mathbf{X}$): four continuous variables (age, income, family size, and years of education) and five binary variables (defined benefit pension status, marital status, two-earner household, IRA participation, and home ownership). All continuous covariates are standardized to the $[0,1]$ interval prior to analysis.
Due to potential unmeasured confounding between 401(k) participation and net financial assets, we focus on estimating the LATE following the approach in Section 5, rather than the overall ATE. Specifically, in this application, the LATE is defined as the effect of 401(k) participation on net financial assets for individuals who would participate in 401(k) if eligible, but not otherwise. The analysis relies on 401(k) eligibility serving as a valid IV, satisfying Assumption 8; for further details, we refer readers to Section 6.2 of Chernozhukov et al. (2018).
We compare eleven estimators in our analysis. First, we include the same five weighting estimators examined in the simulation study in Section 6, namely: (i) Gaussian, (ii) $\ell_{1}$-Laplacian, (iii) $t_{5}$, (iv) Matérn, and (v) energy. Their implementation remains largely unchanged, employing three-way balancing, the median heuristic for hyperparameter tuning, and a regularization parameter of $\lambda=n^{-2}$. The only modification is that the smoothness parameter of the Matérn kernel is set to $s=4.5$ in order to account for the smaller number of continuous covariates. For these five estimators, we report 95% confidence intervals based on both subsampling and bootstrap methods. Second, we consider the same two IPW and CBPS estimators with identical specifications as in the simulation study, reporting 95% bootstrap confidence intervals. Finally, we report four double/debiased machine learning (DDML) estimators, each based on a different nonparametric approach, which are directly available from the online repository of Bach et al. (2022). DDML differs from weighting-based approaches in that it fits flexible nuisance regressions (e.g., treatment and outcome models) and then applies an Neyman orthogonal (influence function-based) estimating equation with cross-fitting to enable valid statistical inference. We include these DDML results as additional benchmarks in this canonical application, even though our simulation study focuses on weighting-based methods.
The results of the data analysis are summarized in Table 2. We first focus on the LATE estimates obtained from our five weighting estimators. The point estimates are largely consistent across different density specifications, with values around \$12,000. Considering the 95% confidence intervals, we further conclude that the LATE is significantly different from zero at the 5% significance level, indicating that 401(k) participation increases net financial assets. Consistent with our simulation findings, the subsampling confidence intervals are wider than the corresponding bootstrap intervals. However, as noted above, the standard bootstrap may fail to achieve the nominal coverage level for our estimator. Therefore, despite being slightly more conservative, we recommend reporting the subsampling confidence intervals as the ultimate measure of uncertainty.
We now turn to a comparison between our method and alternative approaches. Notably, our estimates are closely aligned with those produced by the DDML methods. This agreement with a widely used nonparametric approach further reinforces the credibility of our estimator. In contrast, the point estimates from the IPW and CBPS methods differ substantially, likely due to misspecification in the assumed parametric models. Regarding the confidence intervals, our weighting estimator produces slightly wider intervals (based on subsampling) than the DDML estimator. This modest loss in efficiency may be attributed to two factors: (i) our method does not incorporate outcome information, and (ii) the DDML estimator may attain the semiparametric efficiency bound under mild regularity conditions (Chernozhukov et al., 2018). We anticipate that the augmented weighting estimator discussed in Remark 4 will yield confidence intervals of comparable length. Overall, these empirical findings indicate that our proposed method performs competitively relative to state-of-the-art approaches in this real-world application.
Method Estimate 95% CI 95% CI Length
Distributional Balancing Example 1 Gaussian 11927 SS: (7153, 17140) 9988
Boot: (7405, 14792) 7387
Example 2 $\ell_{1}$-Laplacian 12486 SS: (8984, 16625) 7641
Boot: (8330, 15076) 6746
Example 2 $t_{5}$ 12828 SS: (7813, 17641) 9828
Boot: (8135, 15126) 6990
Example 3 Matérn with $s=4.5$ 12307 SS: (7613, 17234) 9621
Boot: (7887, 15251) 7364
Example 4 Energy 12619 SS: (9399, 16802) 7404
Boot: (8206, 14359) 6153
IPW 444 (-12648, 9684) 22333
CBPS 5769 (-631, 12024) 12655
DDML glmnet 12802 (8999, 16606) 7607
ranger 11792 (8649, 14936) 6287
rpart 12214 (8856, 15573) 6717
xbgoost 11861 (8687, 15035) 6348
Table 2: Summary of the 401(k) dataset analysis. All figures are presented in U.S. dollars.
8 Discussion
In this paper, we introduced a unified nonparametric framework for estimating causal effects using CFD-based balancing weights. In particular, we showed that some existing methods can be viewed as special cases of the proposed framework. Our theoretical results establish the $\sqrt{n}$-consistency of the estimator under suitable regularity conditions. We further highlight several aspects that are often overlooked in practice. First, we clarify the connection between the specification of the density function (or the RKHS kernel) and the corresponding requirements on the outcome regression (see Assumption 4). Second, we provide both theoretical justification and simulation evidence showing that the standard bootstrap may fail, while subsampling offers a valid alternative. We extend our approach to an IV setting to account for potential unmeasured confounding, demonstrating the broad applicability of our method. Simulation and empirical analyses under this IV setting demonstrate that the proposed estimator is robust to model misspecification, in contrast to standard parametric alternatives, and performs competitively with the DDML estimator.
We conclude the paper by outlining potential directions for future research. First, the proposed CFD-based method requires the investigator to specify the density $\omega$ and its associated hyperparameters. We have not addressed how to select these quantities optimally based on the performance of the resulting weighting estimator. Developing a data-driven cross-validation procedure for selecting $\omega$ and its hyperparameters by directly optimizing the causal effect estimator remains an important open direction. Second, the proposed estimator requires solving a constrained QP involving a gram matrix of size $n\times n$. For large $n$, this computation can be burdensome, motivating the development of more efficient algorithms for exceptionally large datasets. We believe that approximation techniques (e.g., Nyström method or sampling-based approaches) are promising directions for future research. Finally, in Section 6, we find that our subsampling confidence intervals are generally conservative. Future research could explore ways to refine subsampling procedures to achieve nominal coverage without being overly conservative, or to develop inferential methods that are theoretically valid for estimators derived from constrained optimization problems.
Supplementary Materials: The supplementary materials contain the proofs of theoretical results, additional numerical results, and R code and data.
Data Availability Statement: The data that support the findings of this study are openly available in DoubleML at https://docs.doubleml.org/stable/index.html.